# Geodesic Slice Sampler for Multimodal Distributions with Strong Curvature

### v1

@@ A.3 The Geodesic Equations
We provide only the necessary background on Riemannian geometry.
A Riemannian metric tensor is the map $g(\cdot,\cdot):T_{\operatorname{\boldsymbol{x}}}\operatorname{\mathcal{M}}\crossproduct T_{\operatorname{\boldsymbol{x}}}\operatorname{\mathcal{M}}\to\operatorname{\mathbb{R}}$ where $T_{\operatorname{\boldsymbol{x}}}\operatorname{\mathcal{M}}$ is the tangent space. The map must satisfy positive definiteness for every $\operatorname{\boldsymbol{x}}\in\operatorname{\mathcal{M}}$. Namely, the components $\operatorname{\boldsymbol{G}}(\operatorname{\boldsymbol{x}})$ are such that for all $\operatorname{\boldsymbol{v}},\boldsymbol{u}\in T_{\operatorname{\boldsymbol{x}}}\operatorname{\mathcal{M}}$, then $g(\operatorname{\boldsymbol{v}},\boldsymbol{u})=\boldsymbol{v}^{\top}\operatorname{\boldsymbol{G}}(\operatorname{\boldsymbol{x}})\boldsymbol{u}>0$ (this is equivalent to saying that $\operatorname{\boldsymbol{G}}(\operatorname{\boldsymbol{x}})$ is positive definite).
Geodesic equations generalize straight lines to curved manifolds. Let $\operatorname{\mathcal{M}}$ be a Riemannian manifold with metric tensor $\operatorname{\boldsymbol{G}}$. The tangent space at a point $\operatorname{\boldsymbol{x}}$ is denoted by $T_{\operatorname{\boldsymbol{x}}}\operatorname{\mathcal{M}}$. A geodesic curve is determined by solving a second-order differential equation with initial conditions $\operatorname{\boldsymbol{x}}_{0}\in\operatorname{\mathcal{M}}$ and $\operatorname{\boldsymbol{v}}_{0}\in T_{\operatorname{\boldsymbol{x}}}\operatorname{\mathcal{M}}$. The metric components are given by $g_{ij}=\operatorname{\boldsymbol{G}}(\operatorname{\boldsymbol{x}})_{ij}$, and their inverse by $g^{km}=\operatorname{\boldsymbol{G}}^{-1}(\operatorname{\boldsymbol{x}})_{km}$. The geodesic equations are:
@@
$\dot{\operatorname{\boldsymbol{v}}}_{k}$ $=-\norm{\operatorname{\boldsymbol{v}}}^{2}_{\Gamma^{k}},\quad\mathrm{for}\ k=1,\ldots,D.$
The Christoffel symbols, using Einstein summation convention, are: $\Gamma^{k}_{ij}=\tfrac{1}{2}g^{km}(\partial_{i}g_{mj}+\partial_{j}g_{im}-\partial_{m}g_{ij})$.
A.4 Sample Uniformly from the Unit Tangent Sphere
Recall the unit tangent sphere is defined as:
@@ A.4 Sample Uniformly from the Unit Tangent Sphere
    Transform $\boldsymbol{v}\leftarrow\boldsymbol{G}^{-\tfrac{1}{2}}(\boldsymbol{x})\boldsymbol{z}$, then $\operatorname{\boldsymbol{v}}$ is a sample from $\mathcal{N}(0,\boldsymbol{G}^{-1}(\boldsymbol{x}))$.
@@
A.5 Hausdorff measure
The volume form of a Riemannian Manifold with metric $\operatorname{\boldsymbol{G}}(\operatorname{\boldsymbol{x}})$ is defined as $V(\operatorname{\textrm{d}\boldsymbol{x}}):=\sqrt{\det\operatorname{\boldsymbol{G}}(\operatorname{\boldsymbol{x}})}\operatorname{\textrm{d}\boldsymbol{x}}$. For the technical details about the volume form interested readers can consult Proposition 2.41 in Lee [2018]. The volume element gives the natural measure on the manifold, analogous to the Lebesgue measure in the Euclidean space [Durmus et al., 2023]. The Hausdorff density is defined as the density which integrates to one with respect to the volume element:
@@ A.5 Hausdorff measure
A.6 Observation and Proposition
@@
Let $p(\operatorname{\boldsymbol{x}})$ be a smooth density function. Let $(\operatorname{\boldsymbol{x}}_{t},\operatorname{\boldsymbol{v}}_{t})$ be the geodesic flow with initial conditions $(\operatorname{\boldsymbol{x}}_{0},\operatorname{\boldsymbol{v}}_{0})$ with respect to the inverse Generative/Monge metric. Then, for $t$ such that $p(\operatorname{\boldsymbol{x}}_{t})\to 0$, the Euclidean velocity norm $\norm{\operatorname{\boldsymbol{v}}_{t}}_{2}$ gets small, implying the low value density regions are covered faster as function of time $t$.
$\operatorname{\boldsymbol{G}}_{Ig}(\operatorname{\boldsymbol{x}})=\left(\frac{p_{0}+\lambda}{p(\operatorname{\boldsymbol{x}})+\lambda}\right)^{2}\operatorname{\boldsymbol{I}},\quad\operatorname{\boldsymbol{G}}_{IM}(\operatorname{\boldsymbol{x}})=\operatorname{\boldsymbol{I}}-\frac{\alpha^{2}}{1+\alpha^{2}\norm{\nabla\ell(\operatorname{\boldsymbol{x}})}^{2}}\nabla\ell(\operatorname{\boldsymbol{x}})\nabla\ell(\operatorname{\boldsymbol{x}})^{\top}.$
Analysis for the Generative Metric
@@
$\norm{\operatorname{\boldsymbol{v}}_{t}}^{2}_{\operatorname{\boldsymbol{G}}_{Ig}}=\norm{\operatorname{\boldsymbol{v}}_{0}}^{2}_{\operatorname{\boldsymbol{G}}_{g}}\quad\forall t.$
@@
$\left(\frac{p_{0}+\lambda}{p(\operatorname{\boldsymbol{x}}_{0})+\lambda}\right)^{2}\norm{\operatorname{\boldsymbol{v}}_{0}}^{2}$ $=\left(\frac{p_{0}+\lambda}{p(\operatorname{\boldsymbol{x}}_{t})+\lambda}\right)^{2}\norm{\operatorname{\boldsymbol{v}}_{t}}_{2}^{2}$
$\left(\frac{p(\operatorname{\boldsymbol{x}}_{t})+\lambda}{p(\operatorname{\boldsymbol{x}}_{0})+\lambda}\right)^{2}\norm{\operatorname{\boldsymbol{v}}_{0}}^{2}$ $=\norm{\operatorname{\boldsymbol{v}}_{t}}_{2}^{2}.$
Since $p(\operatorname{\boldsymbol{x}}_{0})\geq p(\operatorname{\boldsymbol{x}}_{t})$, it follows that $\norm{\operatorname{\boldsymbol{v}}_{0}}_{2}^{2}\geq\norm{\operatorname{\boldsymbol{v}}_{t}}_{2}^{2}$.
@@
Assume a geodesic curve starting at $(\operatorname{\boldsymbol{x}}_{0},\operatorname{\boldsymbol{v}}_{0})$ satisfies $p(\operatorname{\boldsymbol{x}}_{0})\geq p(\operatorname{\boldsymbol{x}}_{t})$ for all $t\geq 0$ and $p(\operatorname{\boldsymbol{x}}_{t})\to 0$. Assume $\operatorname{\boldsymbol{x}}_{0}$ is not a critical point, i.e., $\nabla\log p(\operatorname{\boldsymbol{x}}_{0})\neq 0$. Since $p(\operatorname{\boldsymbol{x}})\geq 0$ is smooth with finite integral, we have $\nabla p(\operatorname{\boldsymbol{x}}_{t})\to 0$. Then:
$\norm{\operatorname{\boldsymbol{v}}_{0}}_{\operatorname{\boldsymbol{G}}_{IM}}^{2}$ $=\norm{\operatorname{\boldsymbol{v}}_{t}}^{2}_{\operatorname{\boldsymbol{G}}_{IM}}$
$\norm{\operatorname{\boldsymbol{v}}_{0}}_{2}^{2}-\frac{\alpha^{2}}{1+\alpha^{2}\norm{\nabla\ell(\operatorname{\boldsymbol{x}}_{0})}^{2}}\left\langle\nabla\ell(\operatorname{\boldsymbol{x}}_{0}),\operatorname{\boldsymbol{v}}_{0}\right\rangle^{2}$ $=\norm{\operatorname{\boldsymbol{v}}_{t}}_{2}^{2}-\frac{\alpha^{2}}{1+\alpha^{2}\norm{\nabla\ell(\operatorname{\boldsymbol{x}}_{t})}^{2}}_{2}\left\langle\nabla\ell(\operatorname{\boldsymbol{x}}_{t}),\operatorname{\boldsymbol{v}}_{t}\right\rangle^{2}$
$\norm{\operatorname{\boldsymbol{v}}_{0}}_{2}^{2}-\frac{\alpha^{2}}{1+\alpha^{2}\norm{\nabla\ell(\operatorname{\boldsymbol{x}}_{0})}^{2}}\left\langle\nabla\ell(\operatorname{\boldsymbol{x}}_{0}),\operatorname{\boldsymbol{v}}_{0}\right\rangle^{2}$ $=\norm{\operatorname{\boldsymbol{v}}_{t}}_{2}^{2}.$
Since $\frac{\alpha^{2}}{1+\alpha^{2}\norm{\nabla\ell(\operatorname{\boldsymbol{x}}_{0})}^{2}}\left\langle\nabla\ell(\operatorname{\boldsymbol{x}}_{0}),\operatorname{\boldsymbol{v}}_{0}\right\rangle^{2}>0$, it follows that $\norm{\operatorname{\boldsymbol{v}}_{0}}_{2}^{2}>\norm{\operatorname{\boldsymbol{v}}_{t}}_{2}^{2}$.
Joint conclusion
Intuitively, this means that if we are on a position $\operatorname{\boldsymbol{x}}_{t}$ along the trajectories we described before where the velocity is of small Euclidean norm $\norm{\operatorname{\boldsymbol{v}}_{t}}_{2}$ then the Riemannian norm of that vector is much higher $\norm{\operatorname{\boldsymbol{v}}_{t}}_{G}$. Then the Exponential map evaluated at that point and velocity is:
$\gamma_{(\operatorname{\boldsymbol{x}},\operatorname{\boldsymbol{v}})}(1)=\gamma_{(\operatorname{\boldsymbol{x}},\operatorname{\boldsymbol{v}}/\norm{\operatorname{\boldsymbol{v}}}_{G})(\norm{\operatorname{\boldsymbol{v}}}_{G})},$ (1)
this means that even a velocity vector with small euclidean norm $\norm{\operatorname{\boldsymbol{v}}_{t}}_{2}$, corresponds to a larger a unit $\norm{\operatorname{\boldsymbol{v}}}_{G}$ of time along the geodesic equation.
Note that for a multimodal distribution in dim $\geq 2$ Observation 1 guarantees that low density regions will be transversed quickly, but we do not have guarantees any that the geodesics given by the inverse Generative and inverse Monge metrics will connect modes.
Note that this analysis also holds for the inverse Monge metric when the geodesic curve approaches a local maximum of the distribution, which explains the behavior observed in Figure 1 around the local maximum values when $\alpha=0.1$.
@@
For a proof of the Proposition in a more general setting, consult 1.5 Proposition of section XII.1 in Amann et al. [2005].
@@ Appendix B Additional Experimental Results
B.1 Unimodal: Logistic regression
Here we denote by $\theta$ the random variable of interest and by $x$ input data. The Logistic regression model [Girolami and Calderhead, 2011]
@@ B.1 Unimodal: Logistic regression
where $\alpha=100$ and $s(\cdot)$ is the Sigmoid function. The Fisher Information Metric plus the Hessian is: $\operatorname{\boldsymbol{G}}(\operatorname{\boldsymbol{x}})=\operatorname{\boldsymbol{X}}^{\top}\operatorname{\boldsymbol{\Lambda}}\operatorname{\boldsymbol{X}}+\alpha^{-1}\operatorname{\boldsymbol{I}}$. Where $\operatorname{\boldsymbol{X}}$ is the covariate matrix and $\operatorname{\boldsymbol{\Lambda}}$ is a diagonal matrix with entries $\operatorname{\boldsymbol{\Lambda}}_{nn}=s(\operatorname{\boldsymbol{x}}_{i}^{\top}\operatorname{\boldsymbol{\theta}})\big{(}1-s(\operatorname{\boldsymbol{x}}_{i}^{\top}\operatorname{\boldsymbol{\theta}})\big{)}$.
References samples are obtained with HMC-NUTS. The Euclidean and Fisher metrics are just as close to the samples, but Fisher and Monge have higher effective sample size and use less shrinkage iterations. In spite of the fact that on flat methods take significantly longer to run. The notation used is $[\mathrm{mean},\mathrm{std}]$ over 5 runs with different seeds.

### v2

@@ A.3 The Geodesic Equations
A Riemannian metric is a smooth, symmetric, and positive-definite tensor $g:T_{\operatorname{\boldsymbol{x}}}\operatorname{\mathcal{M}}\times T_{\operatorname{\boldsymbol{x}}}\operatorname{\mathcal{M}}\to\mathbb{R}$ for each point $\operatorname{\boldsymbol{x}}\in\operatorname{\mathcal{M}}$. In coordinates, the metric is represented by a positive-definite matrix $\operatorname{\boldsymbol{G}}(\operatorname{\boldsymbol{x}})$ such that for all $\operatorname{\boldsymbol{v}},\boldsymbol{u}\in T_{\operatorname{\boldsymbol{x}}}\operatorname{\mathcal{M}}$,
$g(\operatorname{\boldsymbol{v}},\boldsymbol{u})=\operatorname{\boldsymbol{v}}^{\top}\operatorname{\boldsymbol{G}}(\operatorname{\boldsymbol{x}})\boldsymbol{u}.$
Geodesics are curves $\gamma(t)$ on $\operatorname{\mathcal{M}}$ that locally minimize distance and generalize straight lines to curved spaces. They solve the geodesic equation, a second-order ODE determined by the metric. Given initial conditions $\gamma(0)=\operatorname{\boldsymbol{x}}_{0}\in\operatorname{\mathcal{M}}$ and $\dot{\gamma}(0)=\operatorname{\boldsymbol{v}}_{0}\in T_{\operatorname{\boldsymbol{x}}_{0}}\operatorname{\mathcal{M}}$, the geodesic equation in local coordinates is
$\ddot{\gamma}^{k}(t)+\sum_{i,j=1}^{D}\Gamma^{k}_{ij}(\gamma(t))\dot{\gamma}^{i}(t)\dot{\gamma}^{j}(t)=0,\quad\text{for }k=1,\ldots,D,$
where $\Gamma^{k}_{ij}$ are the Christoffel symbols of the second kind, given by
$\Gamma^{k}_{ij}=\tfrac{1}{2}\sum_{m=1}^{D}g^{km}\left(\partial_{i}g_{mj}+\partial_{j}g_{im}-\partial_{m}g_{ij}\right),$
with $g_{ij}=\operatorname{\boldsymbol{G}}(\operatorname{\boldsymbol{x}})_{ij}$ and $g^{km}=(\operatorname{\boldsymbol{G}}^{-1}(\operatorname{\boldsymbol{x}}))_{km}$. Alternatively, defining the position-velocity system with $\operatorname{\boldsymbol{x}}=\gamma(t)$ and $\operatorname{\boldsymbol{v}}=\dot{\gamma}(t)$, the geodesic equations can be expressed as a first-order system (Equation 2):
@@
$\dot{\operatorname{\boldsymbol{v}}}_{k}$ $=-\norm{\operatorname{\boldsymbol{v}}}^{2}_{\Gamma^{k}}\quad\mathrm{for}\ k=1,\ldots,D,$
where $\norm{\operatorname{\boldsymbol{v}}}^{2}_{\Gamma^{k}}=\sum_{i,j=1}^{D}\Gamma^{k}_{ij}(\operatorname{\boldsymbol{x}})\operatorname{\boldsymbol{v}}_{i}\operatorname{\boldsymbol{v}}_{j}$.
A.4 Sampling Uniformly from the Unit Tangent Sphere
Recall the unit tangent sphere is defined by
@@ A.4 Sampling Uniformly from the Unit Tangent Sphere
    Transform $\boldsymbol{v}\leftarrow\boldsymbol{G}^{-\tfrac{1}{2}}(\boldsymbol{x})\boldsymbol{z}$, then $\operatorname{\boldsymbol{v}}$ is distributed according to $\mathcal{N}(0,\boldsymbol{G}^{-1}(\boldsymbol{x}))$.
@@
A.5 The Hausdorff measure
The volume form of a Riemannian manifold with metric $\operatorname{\boldsymbol{G}}(\operatorname{\boldsymbol{x}})$ is defined as $V(\operatorname{\textrm{d}\boldsymbol{x}}):=\sqrt{\det\operatorname{\boldsymbol{G}}(\operatorname{\boldsymbol{x}})}\operatorname{\textrm{d}\boldsymbol{x}}$. For technical details about the volume form, interested readers can consult Proposition 2.41 in Lee [2018]. The volume element gives the natural measure on the manifold, analogous to the Lebesgue measure in Euclidean space [Durmus et al., 2023]. The Hausdorff density is defined as the density which integrates to one with respect to the volume element:
@@ A.5 The Hausdorff measure
An intuitive explanation for the volume element can be thought in terms of change-of-variables in Euclidean space. When transforming coordinates via a diffeomorphism $\phi:\mathbb{R}^{D}\to\mathbb{R}^{D}$, the standard density must be adjusted by the Jacobian determinant to preserve probability mass. That is, $|\det J|$ accounts for local volume distortion.
When $\phi:\mathbb{R}^{D}\to\mathbb{R}^{d}$, where $d>D$ maps from a lower-dimensional Euclidean space onto a manifold embedded in higher dimensions, the Jacobian $J$ is generally rectangular. In this case, the induced Riemannian metric (Pullback metric) on the manifold is $\operatorname{\boldsymbol{G}}(\operatorname{\boldsymbol{x}})=JJ^{\top}$, and the volume change is given by $\sqrt{\det\operatorname{\boldsymbol{G}}(\operatorname{\boldsymbol{x}})}$, which generalizes $|\det J|$.
Thus, the Hausdorff density $p_{\operatorname{\mathcal{H}}}$ adjusts the density $p$ to be properly normalized on the manifold with respect to the intrinsic geometry. This adjustment ensures correct sampling and integration as seen in Proposition 1.
A.6 Observations and Proposition
$\operatorname{\boldsymbol{G}}_{Ig}(\operatorname{\boldsymbol{x}})=\left(\frac{p(\operatorname{\boldsymbol{x}})+\lambda}{p_{0}+\lambda}\right)^{2}\operatorname{\boldsymbol{I}},\quad\operatorname{\boldsymbol{G}}_{IM}(\operatorname{\boldsymbol{x}})=\operatorname{\boldsymbol{I}}-\frac{\alpha^{2}}{1+\alpha^{2}\norm{\nabla\ell(\operatorname{\boldsymbol{x}})}^{2}}\nabla\ell(\operatorname{\boldsymbol{x}})\nabla\ell(\operatorname{\boldsymbol{x}})^{\top}.$
@@ A.6 Observations and Proposition
Let $p(\operatorname{\boldsymbol{x}})$ be a smooth density function. Let $(\operatorname{\boldsymbol{x}}_{t},\operatorname{\boldsymbol{v}}_{t})$ be the geodesic flow with initial conditions $(\operatorname{\boldsymbol{x}}_{0},\operatorname{\boldsymbol{v}}_{0})$ such that $p(x_{0})>0$ with respect to the Inverse Generative metric. Then, for $t$ such that $p(\operatorname{\boldsymbol{x}}_{t})\to 0$ we have $\norm{\operatorname{\boldsymbol{v}}_{t}}_{2}>\norm{\operatorname{\boldsymbol{v}}_{t}}_{0}$.
Analysis for the Inverse Generative metric
@@
$\norm{\operatorname{\boldsymbol{v}}_{t}}^{2}_{\operatorname{\boldsymbol{G}}_{Ig}}=\norm{\operatorname{\boldsymbol{v}}_{0}}^{2}_{\operatorname{\boldsymbol{G}}_{Ig}}\quad\forall t.$
@@
$\left(\frac{p(\operatorname{\boldsymbol{x}}_{0})+\lambda}{p_{0}+\lambda}\right)^{2}\norm{\operatorname{\boldsymbol{v}}_{0}}^{2}$ $=\left(\frac{p(\operatorname{\boldsymbol{x}}_{t})+\lambda}{p_{0}+\lambda}\right)^{2}\norm{\operatorname{\boldsymbol{v}}_{t}}_{2}^{2}$
$\left(\frac{p(\operatorname{\boldsymbol{x}}_{0})+\lambda}{p(\operatorname{\boldsymbol{x}}_{t})+\lambda}\right)^{2}\norm{\operatorname{\boldsymbol{v}}_{0}}^{2}$ $=\norm{\operatorname{\boldsymbol{v}}_{t}}_{2}^{2}.$
Since $p(\operatorname{\boldsymbol{x}}_{0})\geq p(\operatorname{\boldsymbol{x}}_{t})$, it follows that $\norm{\operatorname{\boldsymbol{v}}_{t}}_{2}^{2}\geq\norm{\operatorname{\boldsymbol{v}}_{0}}_{2}^{2}$, and as $p(\operatorname{\boldsymbol{x}}_{t})\to 0$ the quantity is arbitrary large.
Observation 2.
Let $p(\operatorname{\boldsymbol{x}})$ be a smooth density function. Let $(\operatorname{\boldsymbol{x}}_{t},\operatorname{\boldsymbol{v}}_{t})$ be the geodesic flow with initial conditions $(\operatorname{\boldsymbol{x}}_{0},\operatorname{\boldsymbol{v}}_{0})$ with respect to the Inverse Monge metric, such that $\operatorname{\boldsymbol{x}}_{0}$ is a local maximum. Then $\norm{\operatorname{\boldsymbol{v}}_{t}}_{2}\geq\norm{\operatorname{\boldsymbol{v}}_{0}}_{2}$ for all $t\neq 0$.
@@
We consider geodesics emanating from a mode. Let $\operatorname{\boldsymbol{x}}_{0}$ be a mode, meaning that $\nabla\ell(\operatorname{\boldsymbol{x}}_{0})=0$. Consider a geodesic emanating form $\operatorname{\boldsymbol{x}}_{0}$ with velocity $\operatorname{\boldsymbol{v}}_{0}$, it holds
$\norm{\operatorname{\boldsymbol{v}}_{0}}_{2}^{2}-\frac{\alpha^{2}}{1+\alpha^{2}\norm{\nabla\ell(\operatorname{\boldsymbol{x}}_{0})}^{2}}\left\langle\nabla\ell(\operatorname{\boldsymbol{x}}_{0}),\operatorname{\boldsymbol{v}}_{0}\right\rangle^{2}$ $=\norm{\operatorname{\boldsymbol{v}}_{t}}_{2}^{2}-\frac{\alpha^{2}}{1+\alpha^{2}\norm{\nabla\ell(\operatorname{\boldsymbol{x}}_{t})}^{2}}_{2}\left\langle\nabla\ell(\operatorname{\boldsymbol{x}}_{t}),\operatorname{\boldsymbol{v}}_{t}\right\rangle^{2},$
or
$\norm{\operatorname{\boldsymbol{v}}_{0}}_{2}^{2}+\frac{\alpha^{2}}{1+\alpha^{2}\norm{\nabla\ell(\operatorname{\boldsymbol{x}}_{t})}^{2}}_{2}\left\langle\nabla\ell(\operatorname{\boldsymbol{x}}_{t}),\operatorname{\boldsymbol{v}}_{t}\right\rangle^{2}=\norm{\operatorname{\boldsymbol{v}}_{t}}_{2}^{2}.$
We see that $\|\operatorname{\boldsymbol{v}}_{t}\|^{2}\geq\|\operatorname{\boldsymbol{v}}_{0}\|^{2}$, so any geodesic starting from $\operatorname{\boldsymbol{x}}_{0}$ always has a "shrinkage" behavior. So this metric helps bring the entire space close to $\operatorname{\boldsymbol{x}}_{0}$ along geodesics, as it shrinks the space towards (multiple) modes. Also note that the collapsing towards modes depends on how flat the region is and how well-aligned the velocity and the gradient of $\ell$ is. For complicated distributions, the behavior should not depend monotonically on $\alpha$.
Note: For a multimodal distribution of dim $\geq 2$, Observation 1 and Observation 2 guarantee that low-density/increasing-gradient regions the speed increases, but we do not have the guarantee that the geodesics given by the inverse metrics will reach the other modes. The geodesic could twist before reaching the other modes, which could negate the “teleport/move fast" effect.
@@
For a general setting proof of the Proposition consult Section XII.1, Proposition 1.5 in Amann et al. [2005].
@@ Appendix B Additional Experimental Results
B.1 Logistic regression
Here we denote by $\theta$ the random variable of interest and by the $x$ input data. The Logistic regression model [Girolami and Calderhead, 2011] is
@@ B.1 Logistic regression
where $\alpha=100$ and $s(\cdot)$ is the Sigmoid function. The Fisher Information Metric for this probabilistic model including the addition of the the Hessian of the prior is: $\operatorname{\boldsymbol{G}}(\operatorname{\boldsymbol{x}})=\operatorname{\boldsymbol{X}}^{\top}\operatorname{\boldsymbol{\Lambda}}\operatorname{\boldsymbol{X}}+\alpha^{-1}\operatorname{\boldsymbol{I}}$. Where $\operatorname{\boldsymbol{X}}$ is the covariate matrix and $\operatorname{\boldsymbol{\Lambda}}$ is a diagonal matrix with entries $\operatorname{\boldsymbol{\Lambda}}_{nn}=s(\operatorname{\boldsymbol{x}}_{i}^{\top}\operatorname{\boldsymbol{\theta}})\big{(}1-s(\operatorname{\boldsymbol{x}}_{i}^{\top}\operatorname{\boldsymbol{\theta}})\big{)}$.
References samples used for computing the Wasserstein distance are obtained with HMC-NUTS. The samples obtained with the Euclidean and Fisher metrics are just as close to the samples, but Fisher and Monge have higher effective sample size (ESS) and use less shrinkage iterations than the Euclidean metric (See Table 1). The Monge metric uses the parameter $\alpha=1$. $\mathcal{W}$ is the earth mover’s distance. The notation used is $\mathrm{mean}\pm\mathrm{std}$ over 5 runs with different seeds.